If the altitudes of a triangle are in ratio 2:3:4.what are the ratio of their corresponding sides

Grade:10

1 Answers

shaswat roy
13 years ago

the altitudes will be bsinC,asinC,asinB if u wish ,u can take the other values as well using sine rule but ull still get d same ans.

takin ratio of first 2 we get b/a=2/3,ratio of next 2 we get sinC/sinB=3/4 this implies b/c=4/3 using sine rule

thus ratio of the sides are 6:4:3

good luck